The point (4,1) undergoes the following three successive transformation 1 ) Reflection about the y=x-1 2) Translation through a distance 1 units along the positive direction of x-axis 3) Rotation through an angle is 45 about the origin in the anticlockwise direction Then the coordinates of the final point are

Grade:

1 Answers

Arun
5 years ago
Dear student
 
Reflection of (4,1) in x – y – 1 = 0
 
x -4/1 = y-1/-1 = -2(4-1-1)/2
 
x = 4 – 2 = 2
y = 1 + 2 = 3
 
now translation of 1 units in positive direction of x axis
hence coordinates = (3,3)
 
now anticlockwise 45 degree rotation means it will be on y axiss already ot is at 45 degree from origin.
hence coordinates (0, sqrt(18))